OpenStudy (anonymous):

square root of 16 over 81

OpenStudy (ajprincess):

Is it this? \[\sqrt{16/81}\]

OpenStudy (anonymous):

yes

OpenStudy (ajprincess):

\[\sqrt{16}=\sqrt{4*4}=4\]

OpenStudy (ajprincess):

\[\sqrt{81}=\sqrt{9*9}=9\]

OpenStudy (ajprincess):

can u do nw?

OpenStudy (anonymous):

yes thx

OpenStudy (ajprincess):

yw.